4. Using the present value formula find how much you should invest now at 5.9% interest, compounded annually in order to have $10,000, 25 years from now.

Answer by Theo(13342)

you want $10,000 twenty five years from now.
the general formula is f = p * (1 + r) ^ n
f is the future value
p is the present value
r is the interest rate per time period
n is the number of time periods.

your time periods are in year.
the formula uses the rate, not the percent.
your interest rate per year is 5.9% / 100 = .059
your number of years is 25.
your future value is 10,000

the formula becomes 10,000 = p * (1 + .059) ^ 25
solve for p to get p = 10,000 / (1 + .059) ^ 25
solve for p to get p = 2,385.618518

that's your solution.
you would invest 2,385.618518 now in order to have $10,000 twenty five years from now at 5.9% interest rate per year compounded annually.